ABCD is a parallelogram. If AE is equal to 1/2b+1 and EC is equal to b-15, find the measure of AE:
Can you tell me if I have the correct answer, if not, how do you solve this?
32

Answer by nyc_function(2741)

You are almost correct.

The answer is 32 for b.

The distance from A to E = E to C.

1/2b + 1 = b - 15

b = 32

Now that we know b, we must plug that into the equation to find the distance from A to E.

AE = EC

If EC = 32 - 15 or 17, then AE must be 17 also.